# Register Any Point: Scaling 3D Point Cloud Registration by Flow Matching

### v1

@@ 4.3 Ablation Studies
We evaluate our model under different inference and architectural configurations on four testing datasets: 3DMatch and 3DLoMatch for pairwise registration, and ScanNet and map-level KAIST for multi-view registration. The results are summarized in Table 4.
At inference time, we draw $S=5$ random generations and select the one with the smallest rigidity error. Under this default setting [A], our model achieves the highest success rates in three of the four datasets. Compared with [A], the setting without selection [C] generally performs worse, demonstrating the effectiveness of our rigidity-based selection strategy. By comparing [B] and [C], we further observe that enforcing rigidity during sampling consistently improves performance over using the plain Euler integrator. Setting [D] shows that reducing the number of flow integration steps $\kappa$ from 10 to 1 slightly degrades performance but significantly cuts the inference time, as shown in Fig. 6.
The settings [E]-[G] study the impact of model design. Setting [E] uses only the sampled keypoint coordinates as conditions without local descriptors. Its performance drops sharply compared to [A], highlighting the importance of local geometric features. The settings [F] and [G] evaluate the influence of smaller transformer models with blocks $L$ decreasing from 10 to 8 and 6, respectively. As expected, reducing model capacity induces a clear performance trade-off. For additional ablations and results on more datasets, we refer the reader to the supplementary material.
Table 4: Ablation studies. We report the registration success rate (%) on datasets for both pairwise and multi-view registration tasks. Best results are shown in bold, second best are underscored.
@@
3DMatch 3DLoMatch ScanNet KAIST
[A] Ours (S = 5) 95.64 78.22 64.04 87.40
@@
[B] w/o rigidity forcing 94.65 74.63 59.76 83.69
[C] w/o rigidity-based selection 95.27 77.43 63.83 88.12
[D] w/ 1-step generation 90.35 61.05 48.33 52.95
@@
[E] w/o local feature extraction 83.28 50.05 21.82 47.20
[F] w/ $L=6$ transformer blocks 90.90 57.59 45.71 73.77
[G] w/ $L=8$ transformer blocks 93.24 71.63 56.24 85.14
@@ 4.4 Runtime
We further evaluate the inference time of our model on a local workstation with a single NVIDIA A5000 GPU of 24GB VRAM. As shown in Figure 6, the computation time of our model is mainly twofold: the preprocessing time for selecting keypoints and extracting local descriptors and the sampling time of the flow matching. The preprocessing time is shown to increase linearly with the number of input point clouds. In addition, we can sacrifice the accuracy of the registration to reduce the computation time by reducing the number of generation steps $\kappa$, as expected.
Figure 6: Inference time of our model with increasing number of input point clouds NN on a single A5000 GPU. We show the preprocessing time for keypoint selection and local descriptors as well as the sampling time of our flow model on the multi-view 3DMatch dataset.
@@ 5 Conclusion
We presented a generative approach to multi-view point cloud registration that directly generates registered point clouds in a forward pass, bypassing traditional two-stage pipelines. Our method casts registration as conditional generation using rectified point flow matching, where a transformer-based model predicts a continuous velocity field transporting points to registered positions. Holistic reasoning through alternating-attention transformers enables robust registration under low-overlap conditions and effective generalization across diverse scenarios, scales, and sensor modalities. Trained on over 100k samples from 17 diverse datasets, our model achieves state-of-the-art performance on pairwise and multi-view registration benchmarks, with particularly strong results at low overlap ratios. A rigidity-enforcing sampler combined with generation selection further improves test-time performance. Our work advances toward a foundation model for 3D point cloud registration, applicable to SLAM, 3D reconstruction, and robotic manipulation.
Limitations and future work
@@
Our approach assumes scans are recorded in the same scene, like other feed-forward reconstruction methods. Besides, current generation times preclude real-time operation at automotive LiDAR frame rates. By modeling flow in Euclidean space rather than the transformation group, our method could potentially handle non-rigid transformations, though this remains unexplored. Future work may extend to scene flow estimation and merging point maps from photogrammetry and feed-forward 3D reconstruction.
Appendix
In the appendix, we provide the following:
@@
  • Implementation details including point cloud sampling and feature extraction, flow model training, training data curation, testing data details, and evaluation metrics (sections A)
  • Details on the selected baseline methods used in our experiments (section B)
  • Additional experimental results including comprehensive results on all validation and testing datasets as well as extended ablation studies (sections C)
  • Additional qualitative results and failure cases (section D)
Appendix A Implementation Details
Point cloud sampling and feature extraction
@@ Appendix A Implementation Details
For each point cloud $\mathbf{P}_{i}$, we first apply voxel downsampling with voxel size $v_{d}$ to obtain $\mathbf{P}_{i}^{v}$. We then apply statistical outlier removal to $\mathbf{P}_{i}^{v}$ to remove outliers. To ensure uniform sampling density across input point clouds, we determine the sample count $K_{i}$ proportionally to the spatial voxel coverage: we voxelize $\mathbf{P}_{i}^{v}$ with voxel size $v_{c}$ and let $V_{i}$ be the number of remaining points, then set $K_{i}=\lfloor\alpha_{s}V_{i}\rfloor$, where $\alpha_{s}$ is a hyperparameter controlling the FPS sampling density. We apply farthest point sampling (FPS) on $\mathbf{P}_{i}^{v}$ to sample $K_{i}$ points as the feature points $\mathbf{Q}_{i}=\{\mathbf{q}_{i,k}\}_{k=1}^{K_{i}}\in\mathbb{R}^{3\times K_{i}}$. Typically, we got $K_{i}\in[200,2000]$ for our training data. For each feature point $\mathbf{q}_{i,k}\in\mathbf{Q}_{i}$, we extract a local patch by a ball query of radius $r_{s}=20\,v_{d}$ in $\mathbf{P}_{i}^{v}$ and normalize the patch points. We then use the lightweight MiniSpinNet [ao2023cvpr-buffer, seo2025iccv] pretrained on 3DMatch [zeng2017cvpr] as our feature extractor $\mathcal{F}$ to compute a descriptor for each patch with maximum $512$ points and stack them into local features $\mathbf{F}_{i}\in\mathbb{R}^{D\times K_{i}}$, where $D=32$. The pseudocode for this step is shown in Algorithm 1.
@@ 6   Set $K_{i}\leftarrow\lfloor\alpha_{s}V_{i}\rfloor$;
Currently we follow BUFFER-X [seo2025iccv] and use MiniSpinNet as our feature extractor. However, we argue that we can also use other feature extractors to encode local geometry, such as FCGF [choy2019iccv], YOHO [wang2022mm-yoho], and Sonata [wu2025cvpr-sonata].
Flow matching model architecture
We adopt the same flow matching model architecture as in RPF [sun2025neurips] and use a diffusion transformer [peebles2023iccv] with alternating-attention blocks [wang2025cvpr-vggt] for conditional flow matching.
@@
The model’s input consists of the noisy point cloud $\mathbf{X}(t)$ at time step $t$ and a conditioning signal $\mathbf{C}=f_{\text{emb}}(\bar{\mathcal{Q}},\mathcal{F},s)$ obtained via a linear feature embedder $f_{\text{emb}}$. The conditioning $\mathbf{C}$ concatenates three components: (i) local geometric descriptors $\mathcal{F}$ extracted by MiniSpinNet for each sampled keypoint, (ii) positional encodings of the normalized point coordinates $\bar{\mathcal{Q}}$ using a multi-frequency Fourier feature mapping [mildenhall2020eccv], and (iii) a global embedding of the scene scale $s$ to maintain scale awareness, also lifted by the same Fourier feature mapping. The flow matching network $\mathbf{V}_{\theta}$ takes $\mathbf{X}(t)$ and $\mathbf{C}$ as input and predicts the velocity field $\nabla_{t}\mathbf{X}(t)$ that transports the noisy points toward the target registered configuration.
Unlike RPF [sun2025neurips], we do not take the point-wise normals and the scalar view index as additional conditioning signals. We also do not rely on a PTv3-based encoder [wu2024cvpr-ptv3] pretrained for the overlapping prediction task. These design choices make our model simpler and can be applied to various training datasets.
Flow matching model training
@@
We train our model using the Muon Optimizer [liu2025arxiv-muon], with an initial learning rate of $2\times 10^{-3}$ for matrix-like parameters and $2\times 10^{-4}$ for vector-like parameters. From our experiments we find that using Muon instead of AdamW [kingma2015iclr] can achieve faster convergence and better performance. For the learning rate schedule, we halve the learning rate after 400, 550, 700, 850, and 1000 epochs.
We train the model with 8 NVIDIA A100 GPUs with 80GB VRAM each. We train the model in two stages. In the first stage, we train for four days (about 900 epochs) using training samples with point cloud view count $N\in[2,6]$. In the second stage, we train with more views ($N\in[2,24]$) and keep training for three more days (till about 1500 epochs).
Training data curation
@@
The pseudocode for generating the training samples is shown in Algorithm 2.
@@ 30       SaveTrainingSample($\mathcal{M}$);
The details of the training datasets and used parameters for data curation ($N_{\max}$, $F_{\max}$, $d_{\max}$, $\epsilon_{\text{overlap}}$) and point cloud preprocessing ($\alpha_{s}$, $v_{d}$) are shown in Tab. 5. In total, we curate 141k samples containing 520k point clouds, resulting in more than 10 billion points. The data covers a diverse range of scenes across 9 countries on 4 continents, captured by 9 types of LiDAR and 6 types of depth cameras with varying resolutions and scales.
@@
Testing data details
The details of the testing datasets and the evaluation success criteria are shown in Tab. 6.
For pairwise registration, we follow the generalizability benchmark proposed by BUFFER-X [seo2025iccv]. Since BUFFER-X does not directly provide the testing data for ScanNet++ dataset due to the dataset sharing policies, we exclude ScanNet++ from our benchmark. To have enough datasets for indoor depth camera-based point cloud registration evaluation, we additionally add the nothing stands still (NSS) dataset [sun2025jprs]. NSS contains Matterport RGB-D scans of six large indoor areas (offices and construction sites) captured repeatedly over several months as the buildings evolve, leading to large geometric, topological, and appearance changes between fragments [sun2025jprs]. Since our training data do not explicitly consider the temporal change, we adopt the same-stage split in NSS for the evaluation.
For multi-view registration evaluation, except for the commonly used 3DMatch [zeng2017cvpr] and ScanNet [dai2017cvpr] datasets, we select two outdoor LiDAR datasets: Waymo [sun2020cvpr] and KAIST [jung2024ijrr] for the zero-shot multi-view registration evaluation, at both scan- and submap-level. We additionally select two terrestrial laser scanner (TLS) datasets: WHU-TLS [dong2020jprs-whutls] and ETH-TLS [theiler2015jprs] for the zero-shot multi-view registration evaluation. They cover various scenarios such as classroom, park, campus, tunnel, historical building, forest, and mountain, with the scale ranging from 10 m to 1000 m. We only show the qualitative results in the main paper and here we provide the quantitative results in Tab. 9.
Table 6: Summary of the testing datasets used for evaluation. Zero-shot datasets are those not included in the training.
Dataset Scenario Sensor Type # Samples Scale [m] Success Criteria Zero-shot
Pairwise Registration
3DMatch [zeng2017cvpr] USA; office & apartment Kinect, RealSense, etc. Scan 1,623 5 Pointwise RMSE 0.2m No
3DLoMatch [huang2021cvpr-predator] USA; office & apartment Kinect, RealSense, etc. Scan 1,781 5 Pointwise RMSE 0.2m No
NSS [sun2025jprs] USA; office & construction site Matterport camera Scan 1,125 10 TE 0.2m, RE 10 No
TIERS [sier2023rs-tiers] Finland; office & campus Velodyne-16, Ouster-64/128 Scan 870 120 TE 2m, RE 5 Yes
KITTI [geiger2012cvpr] Germany; urban & highway Velodyne-64 Scan 555 160 TE 2m, RE 5 No
Waymo [sun2020cvpr] USA; urban & highway Laser Bear Honeycomb Scan 130 200 TE 2m, RE 5 Yes
KAIST [jung2024ijrr] South Korea; campus Ouster-128, Avia, Aeva Scan 1,991 200 TE 2m, RE 5 Yes
MIT [tian2023iros-kimeramultidata] USA; campus Velodyne-16 Scan 230 160 TE 2m, RE 5 Yes
ETH [pomerleau2012ijrr] Switzerland; park Hokuyo Scan 713 100 TE 2m, RE 5 Yes
NCD [ramezani2020iros] UK; campus Ouster-64 Scan 301 200 TE 2m, RE 5 Yes
Multi-view Registration
3DMatch [zeng2017cvpr] USA; office & apartment Kinect, RealSense, etc. Scan 214 10 TE 0.3m, RE 15 No
ScanNet [dai2017cvpr] USA; office & apartment Structure sensor Scan 155 10 TE 0.3m, RE 15 No
Waymo [sun2020cvpr] USA; urban & highway Laser Bear Honeycomb Scan 77 300 TE 2m, RE 5 Yes
KAIST [jung2024ijrr] South Korea; campus Ouster-128, Avia, Aeva Submap 134 1000 TE 2m, RE 5 Yes
WHU-TLS [dong2020jprs-whutls] China; park, campus, mountain, etc. Riegl, Faro, Leica TLS Scan 11 600 TE 2m, RE 5 Yes
ETH-TLS [theiler2015jprs] Switzerland; courtyard, forest, etc. Z+F, Faro TLS Scan 5 300 TE 2m, RE 5 Yes
Evaluation metrics
We follow previous works on point cloud registration to evaluate the performance of our model using the registration success rate (%) calculated with thresholds on correspondence RMSE for 3DMatch and on translation and rotation error for the other datasets, summarized in Tab. 6. We use threshold settings of previous works [zeng2017cvpr, sun2025jprs, seo2025iccv].
Given the ground truth transformation $\mathbf{T}_{\text{gt}}=[\mathbf{R}_{\text{gt}}|\mathbf{t}_{\text{gt}}]$ and the estimated transformation $\mathbf{T}_{\text{est}}=[\mathbf{R}_{\text{est}}|\mathbf{t}_{\text{est}}]$, we compute the translation error (TE) and rotation error (RE) as:

### v2

@@ 4.4 Ablation Studies
We evaluate our model under different inference and architectural configurations on 3DMatch and 3DLoMatch for pairwise registration, and the cross-domain multi-view registration benchmark for multi-view registration. The results are summarized in Table 3.
Under the default setting [A] with test-time rigidity forcing and $\kappa=10$ step Euler integration, our model achieves the highest success rates across all datasets. Setting [B] removes rigidity forcing during sampling, and setting [C] reduces the number of flow integration steps $\kappa$ from 10 to 1. Both degrade performance, with [B] showing that enforcing rigidity during sampling is crucial and [C] confirming a clear accuracy–speed trade-off with fewer steps. Note that even with a one-step generation, our model outperforms all the baseline methods on the benchmark.
Settings [D]–[F] study the impact of model design. Setting [D] uses only the sampled keypoint coordinates as conditions without local descriptors, and its performance drops sharply compared to [A], especially on the benchmark. Highlighting the importance of local geometric features. Settings [E] and [F] evaluate smaller transformer models with $L$ decreasing from 10 to 8 and 6, respectively. As expected, reducing model capacity induces a clear performance drop. Additional ablations and results are provided in the supplementary material.
Table 3: Ablation studies. We report the registration success rate (%) on datasets for both pairwise and multi-view registration tasks. Best results are shown in bold. On the right figure, we compare the edge success rate of [A]-[C] on the cross-domain multi-view registration benchmark for five different scenario categories.
@@
3DMatch 3DLoMatch Edge Graph
[A] Ours 95.90 78.78 85.92 77.31
@@
[B] w/o rigidity forcing (RF) 94.65 74.63 83.57 72.88
[C] w/ 1-step generation 90.35 61.05 72.87 50.32
@@
[D] w/o local feature extraction 83.28 50.05 43.07 26.85
[E] w/ $L=8$ transformer blocks 93.24 71.63 81.08 72.19
[F] w/ $L=6$ transformer blocks 90.90 57.59 65.64 54.35
@@ 4.5 Runtime
We further evaluate the inference time of our model on a local workstation with a single NVIDIA A5000 GPU of 24 GB VRAM. As shown in Fig.˜5 (left), at two views, RAP is slightly slower than KISS-Matcher and BUFFER-X due to the flow-matching inference overhead. As the number of views grows, however, the baselines scale poorly due to the cost of all-pairs pairwise registration and pose graph optimization, while RAP’s runtime remains nearly linear, quickly becoming the fastest option overall. The right plot breaks down RAP’s compute into two components: feature extraction and flow-matching inference. Feature extraction dominates at low token counts but scales linearly with the total number of feature point tokens, while flow-matching inference grows more steeply due to the quadratic complexity of global attention. Overall, the runtime scales gracefully, and the number of Euler steps $\kappa$ can provide a direct accuracy–speed trade-off at flow matching generation time.
Figure 5: Multi-view registration run time analysis on the cross-domain multi-view registration benchmark. Left: Mean runtime per sample vs. view count for RAP, BUFFER-X*+PGO, and KISS-Matcher+PGO. Right: RAP runtime vs. total feature point token count fed into the transformer, split into feature extraction backbone and flow-matching inference time. Runtimes reported on a single Nvidia A5000 GPU excluding IO.
@@ 5 Conclusion
We presented RAP, a generative flow-matching model for single-stage multi-view point cloud registration. By casting registration as conditional point flow generation with an alternating-attention transformer, RAP achieves holistic multi-view reasoning without a pairwise stage or pose graph optimization, yielding strong robustness under low overlap. Trained on over 100k multi-view registration instances from 17 diverse datasets and evaluated zero-shot on our proposed cross-domain benchmark, RAP achieves state-of-the-art performance across object, indoor, outdoor, TLS, and map domains. A rigidity-enforcing sampler provides further test-time gains. We hope RAP serves as a step toward a universal registration foundation model for applications such as SLAM, 3D reconstruction, and robotic manipulation.
@@
Limitations and future work  Our approach assumes that the scans to be registered are recorded in the same environment, like other feed-forward reconstruction methods [wang2025cvpr-vggt, wang2025arxiv-pi3]. By modeling flow in Euclidean space rather than the transformation group, our method could potentially handle non-rigid transformations, though this remains unexplored. Future work may extend to scene flow estimation [huang2022eccv] and merging point maps from photogrammetry [schoenberger2016cvpr-sfm] and feed-forward 3D reconstruction [wang2025cvpr-vggt, wang2025arxiv-pi3].
Supplementary Material
Appendix A Overview
In the supplementary material, we provide the following:
@@ Appendix A Overview
  • Implementation details including point cloud sampling and feature extraction, flow model architecture and training, training data curation, pairwise registration testing data, and evaluation metrics (Section B)
  • Details on the selected baseline methods used in our experiments (Section C)
  • Additional experimental results including pairwise registration with low overlap, indoor multi-view registration on 3DMatch and ScanNet, cross-domain multi-view registration benchmark, and offline SLAM pose estimation (Section D)
  • Additional information of the cross-domain multi-view registration benchmark (Section E)
  • Additional qualitative results and failure cases (Section F)
Appendix B Implementation Details
B.1 Point cloud sampling and feature extraction
@@
For each point cloud $\mathbf{P}_{i}$, we first apply voxel downsampling with voxel size $v_{d}$ to obtain $\mathbf{P}_{i}^{v}$. We then apply statistical outlier removal to $\mathbf{P}_{i}^{v}$ to remove outliers. To ensure uniform sampling density across input point clouds, we determine the sample count $K_{i}$ proportionally to the spatial voxel coverage: we voxelize $\mathbf{P}_{i}^{v}$ with voxel size $v_{c}$ and let $V_{i}$ be the number of remaining points, then set $K_{i}=\lfloor\alpha_{s}V_{i}\rfloor$, where $\alpha_{s}$ is a hyperparameter controlling the FPS sampling density. We apply farthest point sampling (FPS) on $\mathbf{P}_{i}^{v}$ to sample $K_{i}$ points as the feature points $\mathbf{Q}_{i}=\{\mathbf{q}_{i,k}\}_{k=1}^{K_{i}}\in\mathbb{R}^{3\times K_{i}}$. Typically, we got $K_{i}\in[200,5000]$ for our training data. For each feature point $\mathbf{q}_{i,k}\in\mathbf{Q}_{i}$, we extract a local patch by a ball query of radius $r_{s}=20\,v_{d}$ in $\mathbf{P}_{i}^{v}$ and normalize the patch points. We then use the lightweight MiniSpinNet [ao2023cvpr-buffer, seo2025iccv] pretrained on 3DMatch [zeng2017cvpr] as our feature extractor $\mathcal{F}$ to compute a descriptor for each patch with maximum $512$ points and stack them into local features $\mathbf{F}_{i}\in\mathbb{R}^{D\times K_{i}}$, where $D=32$. During inference, there are only two tunable hyperparameters: the sampling density ratio $\alpha_{s}$ and the downsampling voxel size $v_{d}$. For ease of use, we set $\alpha_{s}=0.2$ by default and determine $v_{d}$ adaptively based on the scene scale, requiring no scene-specific hyperparameter tuning.
The pseudocode for this step is shown in Algorithm˜1.
@@ 6   Set $K_{i}\leftarrow\lfloor\alpha_{s}V_{i}\rfloor$;
Currently, we use MiniSpinNet [ao2023cvpr-buffer] as our feature extractor. However, our framework is compatible with other feature extractors for encoding local geometry, including task-specific ones trained for registration (e.g., FCGF [choy2019iccv] and YOHO [wang2022mm-yoho]) as well as more general self-supervised backbones based on point transformers [wu2024cvpr-ptv3] (e.g., Sonata [wu2025cvpr-sonata] and Utonia [zhang2026arxiv-utonia]). We leave the exploration of other feature extractors as our future work.
B.2 Flow matching model architecture
Following RPF [sun2025neurips], we use a diffusion transformer [peebles2023iccv] with alternating-attention blocks [wang2025cvpr-vggt] for conditional flow matching.
@@ B.2 Flow matching model architecture
The model’s input consists of the noisy point cloud $\mathbf{X}(t)$ at time step $t$ and a conditioning signal $\mathbf{C}=f_{\text{emb}}(\bar{\mathcal{Q}},\mathcal{F})$ obtained via a linear feature embedder $f_{\text{emb}}$. The conditioning $\mathbf{C}$ concatenates three components: (i) local geometric descriptors $\mathcal{F}$ extracted by MiniSpinNet for each sampled keypoint, (ii) positional encodings of the normalized point coordinates $\bar{\mathcal{Q}}$ using a multi-frequency Fourier feature mapping [mildenhall2020eccv]. The flow matching network $\mathbf{V}_{\theta}$ takes $\mathbf{X}(t)$ and $\mathbf{C}$ as input and predicts the velocity field $\nabla_{t}\mathbf{X}(t)$ that transports the noisy points toward the target registered configuration.
Unlike RPF [sun2025neurips], we do not take the point-wise normals and the scalar view index as additional conditioning signals. We also do not rely on a PTv3-based encoder [wu2024cvpr-ptv3] pretrained for the overlapping prediction task. These design choices make our model simpler and can be applied to various cross-domain training datasets.
B.3 Flow matching model training
@@
We train our model using the Muon Optimizer [liu2025arxiv-muon], with an initial learning rate of $5\times 10^{-3}$ for matrix-like parameters and $5\times 10^{-4}$ for vector-like parameters. From our experiments we find that using Muon instead of AdamW [kingma2015iclr] can achieve faster convergence and better performance. For the learning rate schedule, we halve the learning rate after 200, 240, 280, 320, 360, 400, and 500 epochs. We train the model for three days with about 120k iterations using 32 NVIDIA A100 GPUs with 80 GB VRAM each. The training data consists of registration instances with view counts ranging from 2 to 16.
B.4 Training data curation
@@
The pseudocode for generating the training samples is shown in Algorithm˜2.
@@ 30       SaveTrainingSample($\mathcal{M}$);
The details of the training datasets and used parameters for data curation ($N_{\max}$, $F_{\max}$, $d_{\max}$, $\epsilon_{\text{overlap}}$) and point cloud preprocessing ($\alpha_{s}$, $v_{d}$) are shown in Tab.˜4. In total, we curate 141k samples containing 520k point clouds, resulting in more than 10 billion points. The data covers a diverse range of scenes across 9 countries on 4 continents, captured by 9 types of LiDAR and 6 types of depth cameras with varying resolutions and scales.
@@
B.5 Pairwise registration testing data details
The details of the six adopted testing datasets for pairwise registration evaluation and the evaluation success criteria are shown in Tab.˜5.
Table 5: Summary of the testing datasets used for pairwise registration evaluation.
Dataset Scenario Sensor Type # Samples Scale [m] Success Criteria
ModelNet [wu2015cvpr] Synthetic; object CAD Object 1,266 1 TE 0.1m, RE 5
3DMatch [zeng2017cvpr] USA; office & apartment Kinect, RealSense, etc. Scan 1,623 5 Pointwise RMSE 0.2m
3DLoMatch [huang2021cvpr-predator] USA; office & apartment Kinect, RealSense, etc. Scan 1,781 5 Pointwise RMSE 0.2m
NSS [sun2025jprs] USA; office & construction site Matterport camera Scan 1,125 10 TE 0.2m, RE 10
ETH [pomerleau2012ijrr] Switzerland; park Hokuyo Scan 713 100 TE 2m, RE 5
KITTI [geiger2012cvpr] Germany; urban & highway Velodyne-64 Scan 555 160 TE 2m, RE 5
B.6 Evaluation metrics
We evaluate pairwise registration performance using the registration success rate (%), computed with thresholds on correspondence RMSE for 3DMatch and on translation and rotation errors for all other datasets; the exact thresholds are summarized in Tab.˜5 and follow the settings of prior works [zeng2017cvpr, sun2025jprs, seo2025iccv].
Given the ground-truth transformation $\mathbf{T}_{\text{gt}}=[\mathbf{R}_{\text{gt}}\mid\mathbf{t}_{\text{gt}}]$ and the estimated transformation $\mathbf{T}_{\text{est}}=[\mathbf{R}_{\text{est}}\mid\mathbf{t}_{\text{est}}]$, the translation error (TE) and rotation error (RE) are defined as: